# DF-Calib: Targetless LiDAR-Camera Calibration via Depth Flow

### v1

@@ III-B Depth Generation
Inspired by the concept of optical flow estimation, we introduce depth flow $\mathbf{f}\in\mathbb{R}^{2\times W\times H}$ as a pixel-wise displacement field between two depth maps for LiDAR-camera extrinsic calibration, capturing both horizontal and vertical shifts. The depth flow represents dense pixel correspondences between two depth maps while inherently establishes 2D-3D correspondences between the camera and the LiDAR data, which are subsequently used for extrinsic estimation.
@@ III-C Calibration Network
Our network architecture is inspired by SEA-RAFT [31] and integrates LiDAR-projected and camera-estimated depth information to implement depth flow estimation.
III-C1 Feature Encoder
As introduced in sect. III-B, we complete $D_{L}^{s}$ to obtain $D_{L}^{d}$, which enhances the feature similarity between LiDAR and the camera depth maps. While $D_{L}^{s}$ contains sparse, discontinuous data with limited point features, $D_{L}^{d}$ incorporates more expressive line and surface features, making it texturally closer to $D_{I}^{d}$, as shown in Fig. 2. As a result, we can employ a shared depth feature encoder for both depth map $D_{L}^{d}$ and $D_{I}^{d}$ after they are normalized to the range $[-1,1]$, reducing network complexity and facilitating the extraction of consistent depth-to-depth features, thereby enhancing calibration performance. Additionally, $D_{L}^{d}$ and $D_{I}^{d}$ are fed into a single context feature encoder to extract context information $C$ from both LiDAR and camera data, which is subsequently decomposed into context features $F_{C}$ and the initial depth flow $\mathbf{f}_{init}$.
@@ III-C1 Feature Encoder
However, depth completion introduces noise and blur effects, which may lead to the extraction of erroneous features, potentially undermining the accuracy of the extrinsic calibration. To address this issue, we utilize a depth map encoder to generate a reliability map $R$ from $D_{L}^{s}$, guiding the model to prioritize features extracted from valid pixels. In the subsequent stage, $R$ is used to enhance the accuracy of feature correlation.
III-C2 Iterative Refinement
Based on the features extracted in the previous stage, we first apply element-wise multiplication between the features $F_{L}$ extracted from $D_{L}^{d}$ and the reliability map $R$ to obtain the refined features $F_{L}^{R}$. Subsequently, following SEA-RAFT [31], we construct a multi-scale feature correlation pyramid $F_{multi}$ between the depth features $F_{I}$ extracted from $D_{I}^{d}$ and $F_{L}^{R}$ by computing the dot product between corresponding pixel pairs at each scale. The refinement block takes context features $F_{C}$, multi-scale features $F_{multi}$, and the depth flow $\mathbf{f}$ as inputs, where $\mathbf{f}$ is initially set to the initial depth flow $\mathbf{f}_{init}$. It then iteratively updates $\mathbf{f}$ through a recurrent refinement process, progressively enhancing the accuracy of the estimated depth flow. The final refined depth flow, representing the 2D-3D correspondences, is subsequently processed by PnP with RANSAC to compute the extrinsic parameter $\mathbf{T}_{est}$.
III-D Loss Function
@@
Since the ground truth depth flow $\mathbf{f}_{gt}$ is derived from the projection of the sparse LiDAR point cloud, it only contains valid information at a sparse set of pixels. Directly applying a standard $\mathcal{L}_{1}$ loss to all pixels would allow these invalid pixels to negatively impact the estimation performance of the final depth flow. To mitigate this issue, we introduce a valid mask $v(p)$ that filters out invalid pixels, defined as follows:
@@
During depth flow estimation, regions characterized by significant depth variations generally correspond to object boundaries or other critical structural features, which are paramount for accurate flow estimation. Conventional $\mathcal{L}_{1}$ loss function treats all pixels equally, which leads to suboptimal learning performance, particularly in areas with sharp depth discontinuities, where flow estimation errors are more perceptible and visually significant. Emphasizing these regions helps prioritize critical features, thereby enhancing depth flow estimation accuracy.
Building upon this principle, we propose the perceptually weighted sparse flow (PWSF) loss, which adaptively emphasizes depth flow errors in these critical regions while maintaining overall sparsity within the flow field. Specifically, the PWSF loss is formulated as follows:
$\begin{split}\mathcal{L}_{PWSF}=&\frac{1}{S}\sum_{p}\Big{(}\alpha_{i}(p)\alpha_{j}(p)\left\|\mathbf{f}_{gt}(p)-\mathbf{f}_{pre}(p)\right\|_{1}\\ &+\left(1-\alpha_{i}(p)\right)\left\|\mathbf{f}_{gt}(p)-\mathbf{f}_{pre}(p)\right\|_{1}\Big{)}v(p),\end{split}$ (4)
where $S=\sum_{p}v(p)$ is the sum of the valid pixels. $\alpha\in\mathbb{R}^{2\times H\times W}$ represents the model-predicted importance map, $\alpha_{i}\in\mathbb{R}^{1\times H\times W}$ representing the likelihood of each pixel belonging to the salient regions, indicating its importance, and $\alpha_{j}\in\mathbb{R}^{1\times H\times W}$ controlling the weighting of the corresponding loss term. Based on the predicted importance of each pixel, the importance map adjusts the weight of the loss components. For pixels with higher predicted importance, indicating they are likely to lie within salient regions, the weighted $\mathcal{L}_{1}$ loss is emphasized. In contrast, the standard $\mathcal{L}_{1}$ loss dominates. This strategy ensures the overall sparsity of the flow field while effectively prioritizing accuracy in regions with critical structural details.
@@
where $N$ denotes the total number of iterations, and $\gamma<1$ is an exponential decay factor that reduces the influences of later iterations. The term $\mathcal{L}_{PWSF}^{i}$ is the PWSF loss at the $i$-th iteration. This weighted summation ensures that earlier iterations contribute more significantly to the optimization process, promoting a more stable and progressive refinement of the depth flow.
IV Experiments
TABLE I: Different environmental setups
@@ IV Experiments
TABLE II: Comparison results on different environmental setups
[…]
@@
DF-Calib 5.598 5.697 4.553 6.544 0.326 0.388 0.325 0.266
In this section, we conduct comprehensive experiments on the KITTI Raw [34], KITTI Odometry [34], and KITTI-360 [35] datasets to compare DF-Calib with the state-of-the-art approaches, validating its effectiveness and generalization capability.
IV-A Dataset Preparation
@@
Since there is no standardized experimental setup for LiDAR-camera extrinsic calibration, different experimental configurations are set for comparison, as shown in Table I.
IV-A1
@@
IV-A2
Exp2 follows the setup of CalibFormer [12], where sequences 01 to 21 of the KITTI Odometry dataset are used for training and validation, while sequence 00 is used for testing. The initial perturbation is set to $(10^{\circ},0.25\mathrm{m})$.
IV-A3
@@
IV-B Experimental Settings
IV-B1 Implementation Details
All images are cropped to a resolution of 960 $\times$ 320, with the camera intrinsic parameter adjusted accordingly for LiDAR point cloud projection. The AdamW optimizer is used with a weight decay of $1\times 10^{-5}$ and an initial learning rate of $1\times 10^{-4}$. The OneCycleLR scheduler is employed for learning rate adjustment. The model is trained for 100 epochs with a batch size of 6, and the number of iterative refinement is set to 4 during training. All the experiments are conducted on a single NVIDIA RTX 4090 GPU.
IV-B2 Evaluation Metrics
@@
IV-C Experimental Results
@@
In this section, we analyze the quantitative and qualitative results of DF-Calib compared to other methods. Additionally, we examine its runtime performance.
@@
Quantitative Results. As shown in Table II, DF-Calib is evaluated against seven state-of-the-art approaches across various initial deviations and datasets, including NetCalib [32], DXQ-Net [15], CALNet [27], PSNet [33], CalibFormer [12], CalibDepth [16] and CalibLG [13].
IV-C1 Exp1
@@
DF-Calib achieves a mean translation error of $0.635\mathrm{cm}$ and a mean rotation error of $0.045^{\circ}$. When compared to DXQ-Net, a matching-based method without region prioritization, DF-Calib maintains comparable rotation error while greatly reducing the translation error by $18.0\%$. This improvement stems from the PWSF loss, which prioritizes salient regions in LiDAR and camera data, enhancing depth flow estimation and calibration precision.
IV-C2 Exp2&\&&3
@@
DF-Calib achieves a mean translation error of $1.065\mathrm{cm}$ and a mean rotation error of $0.058^{\circ}$ in Exp2, surpassing CalibFormer, which regresses the extrinsic from a camera image and a sparse LiDAR depth map, by $10.4\%$ and $58.9\%$, respectively. In Exp3, it attains a mean translation error of $2.271\mathrm{cm}$ and a mean rotation error of $0.091^{\circ}$, outperforming CalibLG, another regression-based method using the same inputs, by $40.4\%$ and $62.1\%$. These improvements arise from DF-Calib’s reformulation of the calibration task as an intra-modality depth flow estimation problem, enhancing feature consistency and significantly boosting the accuracy of both translation and rotation estimation.
IV-C3 Exp4
@@
DF-Calib achieves a mean translation error of $5.598\mathrm{cm}$ and a mean rotation error of $0.326^{\circ}$ on the KITTI-360 dataset without being trained on it, outperforming CalibDepth, the only open-source method among the baselines, by $70.2\%$ and $86.5\%$, respectively. This strong generalization ability stems from DF-Calib being a matching-based method, which decouples calibration from camera intrinsics and enhances adaptability to diverse camera configurations.
Figure 3: Overlay of RGB images with LiDAR projections using uncalibrated, calibrated, and GT extrinsics, respectively, across KITTI Odometry, KITTI Raw and KITTI-360 datasets.
@@
Qualitative Results. As shown in Fig. 3, DF-Calib’s estimated extrinsics closely align with the ground truth across diverse scenarios and initial deviation conditions, highlighting its robustness to a wide range of initial errors and effectiveness in handling complex environments.
Runtime Analysis. We optimize MoGe by removing unnecessary operations, and reducing depth estimation to 38 ms. In parallel, LiDAR projection and depth completion take 12 ms. Then depth flow estimation takes 31 ms, and the PnP solver 25 ms, totaling about 95 ms. This enables calibration at about 11 Hz, acceptable for real-time applications.
TABLE III: Ablation studies on different components in DF-Calib
@@
[…]
@@
  • D, S, R, and P represent Completed Dense Depth Map, Shared Feature Encoder, Reliability Map, and PWSF Loss, respectively.
IV-D Ablation Study

### v2

@@ 3.2 Depth Generation
Inspired by optical flow, we define depth flow $\mathbf{f}\in\mathbb{R}^{2\times W\times H}$ as a pixel-wise displacement between two depth maps, capturing horizontal and vertical shifts. It encodes dense pixel correspondences and establishes 2D-3D correspondences between the camera and LiDAR for extrinsic estimation.
@@ 3.3 Probabilistic Calibration Network
Our network architecture, inspired by SEA-RAFT [29], integrates LiDAR-projected and camera-predicted depth maps into a unified probabilistic framework for depth flow inference.
3.3.1 Depth Feature Encoder
As $D_{L}^{d}$ incorporates richer geometric structures, it becomes more texturally similar to the dense camera depth map $D_{I}^{d}$, as shown in Fig. 3. To exploit this enhanced similarity, we normalize both depth maps to the range [1,1][-1,1] and feed them into a unified ResNet-34 depth encoder. This enables consistent, modality-invariant feature extraction, improves training efficiency through parameter sharing, and helps the network learn a shared feature space, improving cross-modal calibration accuracy. Additionally, a separate ResNet-34 context encoder extracts context features $F_{C}$ and the initial flow $\mathbf{f}_{init}$ from the normalized depth maps
@@ 3.3.1 Depth Feature Encoder
However, depth completion introduces noise and blur, leading to unreliable features and degraded calibration accuracy. To address this issue, we predict a reliability map $R$ from the sparse depth $D_{L}^{s}$ via a depth encoder, guiding feature extraction by emphasizing reliable pixels. Specifically, we perform element-wise multiplication between the features $F_{L}$ extracted from the completed depth $D_{L}^{d}$ and $R$ to obtain reliability-weighted features $F_{L}^{R}$, which enhances the accuracy of feature correlation.
Following RAFT [26], we construct a multi-scale 4D correlation volume by computing all-pairs dot products between the depth features $F_{I}$ extracted from $D_{I}^{d}$ and the downsampled $F_{L}^{R}$ at each level $k$:
$V_{k}=F_{I}\circ\texttt{AvgPool}(F_{L}^{R},2^{k})^{\top}\in\mathbb{R}^{h\times w\times\frac{h}{2^{k}}\times\frac{w}{2^{k}}},$ (3)
where $\circ$ denotes the dot product between feature vectors. We set $(h,w)=\frac{1}{8}(H,W)$ and use $k=4$ to balance efficiency and accuracy. The resulting correlation volume provides rich similarity cues across scales, serving as the input to the subsequent probabilistic flow decoder.
3.3.2 Probabilistic Flow Decoder
To address the ambiguity in depth flow estimation caused by occlusions and dynamic objects, we design a probabilistic decoder that jointly predicts the depth flow and uncertainty, allowing the model to express confidence and improve robustness in challenging regions. Given a pair of depth maps $X=\left(D_{I}^{d},D_{L}^{d}\right)$, the goal is to estimate a flow $\mathbf{f}\in\mathbb{R}^{2\times W\times H}$ that warps $D_{L}^{d}$ to align with $D_{I}^{d}$. Instead of directly regressing $\mathbf{f}$ via a network $F$ with parameters $\theta$, we model its conditional probability distribution $p(\mathbf{f}|X;\theta)$. To achieve this, the decoder predicts the parameters $\Phi(X;\theta)$ of a family of distributions $p(\mathbf{f}|X;\theta)=p(\mathbf{f}|\Phi(X;\theta))=\prod_{uv}p(f_{uv}|\varphi_{uv}(X;\theta))$. Here, $f_{uv}\in\mathbb{R}^{2}$ is the flow at pixel $(u,v)$, while $\varphi_{uv}\in\mathbb{R}^{n}$ are the corresponding predicted parameters. For simplicity, we omit the subscript $uv$ in what follows. In this case, the density is represented using Laplace distributions as follows:
$p(f|\mu,\sigma^{2})=\frac{1}{\sqrt{2\sigma_{u}^{2}}}e^{-\sqrt{\frac{2}{\sigma_{u}^{2}}}|u-\mu_{u}|}.\frac{1}{\sqrt{2\sigma_{v}^{2}}}e^{-\sqrt{\frac{2}{\sigma_{v}^{2}}}|v-\mu_{v}|},$ (4)
where $f=(u,v)\in\mathbb{R}^{2}$ denotes the flow vector with conditionally independent components. The network predicts the mean $\mu=[\mu_{u},\mu_{v}]^{T}$ as the flow and a shared variance $\sigma^{2}$ to represent uncertainty, such that $\sigma_{u}^{2}=\sigma_{v}^{2}=\sigma^{2}$ for efficiency and stability.
The decoder follows an iterative refinement strategy inspired by RAFT [26], but differs in that our output is four-dimensional. Specifically, it predicts distribution parameters $\Phi$ consisting of a two-dimensional mean $\boldsymbol{\mu}$ and a one-dimensional variance $\boldsymbol{\sigma}^{2}$, as well as a one-dimensional outlier probability map $\boldsymbol{\alpha}$. The update block takes the current flow estimate, the 4D correlation volume, and context features $F_{C}$ as input at each iteration. The final optimized mean $\boldsymbol{\mu}$ is treated as the predicted flow and passed to a PnP solver to estimate the extrinsics. Both the predicted variance $\boldsymbol{\sigma}^{2}$ and outlier probability $\boldsymbol{\alpha}$ are used in the loss function to supervise the distribution, as detailed in Sec. 3.4.
3.4 Training Objective
3.4.1 Valid Mask
@@
As $\mathbf{f}_{gt}$ is derived from the sparse LiDAR projection, it is only valid at limited pixels. Applying loss over all pixels would introduce noise from invalid regions and degrade depth flow estimation. To mitigate this, we introduce a valid mask $v(p)$ to filter out invalid pixels, defined as:
@@
3.4.2 Negative Log-likelihood Loss
Previous probabilistic regression methods [2, 24] often employ the negative log-likelihood (NLL) loss, which, for the Laplace distribution, is defined as
$\mathcal{L}(\boldsymbol{\mu_{gt}},\boldsymbol{\mu},\boldsymbol{\sigma}^{2})=$ $\frac{1}{HW}\sum_{p}\Bigg{[}\log\left(\frac{1}{2\sigma^{2}(p)}\right)$
$+\sqrt{\frac{2}{\sigma^{2}(p)}}\left\|\boldsymbol{\mu_{gt}}(p)-\boldsymbol{\mu}(p)\right\|_{1}\Bigg{]},$ (6)
where $\boldsymbol{\mu}$ and $\boldsymbol{\mu}_{gt}$ are the predicted and ground truth flows, and $\boldsymbol{\sigma}^{2}$ is the predicted per-pixel variance, which modulates confidence by downweighting high-uncertainty errors. However, this formulation is not well-suited for flow estimation tasks[36], especially depth flow, as it exhibits two inherent limitations: the application of uniform supervision irrespective of prediction confidence, and the tendency to overfit uncertain regions by inflating the predicted variance.
3.4.3 Perceptually Weighted Sparse Flow Loss
To address this, we propose the Perceptually Weighted Sparse Flow (PWSF) loss, which leverages the model’s estimated confidence to adaptively suppress the influence of outliers during supervision. Specifically, the loss is weighted by a predicted outlier probability map $\boldsymbol{\alpha}\in[0,1]^{H\times W}$, which down-weights supervision in regions likely to contain mismatches or occlusions. The PWSF loss comprises two complementary components, defined as:
$\mathcal{L}_{\mathrm{PWSF}}=\frac{1}{N_{valid}}\sum_{p}v(p)\Big{[}$ $(1-\alpha(p))\,\mathcal{L}(\mathbf{f}_{gt},\boldsymbol{\mu},\mathbf{2})$
$+\alpha(p)\,\mathcal{L}(\mathbf{f}_{gt},\boldsymbol{\mu},\boldsymbol{\sigma}^{2})\Big{]},\vskip-10.00002pt$ (7)
where $N_{valid}=\sum_{p}v(p)$ denotes the total number of valid pixels. When $\alpha(p)$ is small, indicating a high likelihood of inliers, the supervision is dominated by a standard L1 loss, which corresponds to the NLL loss of a Laplace distribution with fixed variance $\boldsymbol{\sigma}^{2}=\mathbf{2}$. Conversely, when $\alpha(p)$ is large, indicating likely outliers, a variance-weighted formulation takes precedence, allowing the network to attenuate the influence of unreliable predictions via a learned uncertainty map.
The PWSF loss adaptively downweights unreliable predictions while preserving strong supervision for confident ones. It mitigates variance inflation in unreliable regions and maintains overall flow sparsity, leading to improved robustness in real-world scenarios.
3.4.4 Total Loss
@@
where $N$ denotes the total number of iterations, and $\gamma=0.8$ is an exponential decay factor that reduces the influence of later iterations. $\mathcal{L}_{PWSF}^{i}$ is the PWSF loss at the $i$-th iteration. This weighted summation ensures that earlier iterations contribute more to the optimization process, promoting a more stable and progressive refinement.
Table 1: Different experimental configurations
@@
4 Experiments
Table 2: Comparison results on different experimental configurations
Dataset Method Translation ($\mathrm{cm}$) $\downarrow$ Rotation ($\mathrm{\SIUnitSymbolDegree}$) $\downarrow$
@@ 4 Experiments
UniCalib 9.664 10.408 9.299 9.285 0.381 0.443 0.386 0.314
In this section, we evaluate UniCalib on the KITTI Raw [5], KITTI Odometry [5], and KITTI-360 [15] datasets to compare it with state-of-the-art approaches, validating its effectiveness and generalization capability.
4.1 Experimental Setup
4.1.1 Dataset Preparation
@@
As no standardized setting exists for LiDAR-camera extrinsic calibration, we establish several configurations to facilitate fair comparisons, as presented in Tab. 1.
@@
Exp2 follows the configuration of CalibFormer [32], where sequences 01 to 21 of the KITTI Odometry dataset are used for training and validation, while sequence 00 is used for testing. The initial perturbation is set to $(10^{\circ},0.25\mathrm{m})$.
@@
4.1.2 Implementation Details
Our network is adapted from SEA-RAFT [29], with a modified ResNet-34 backbone for depth feature extraction. The probabilistic flow decoder is implemented as described in Sec. 3.3.2. Data augmentation includes random cropping and horizontal flipping. All images are cropped to a resolution of 960 $\times$ 320, and the camera intrinsics are adjusted accordingly to ensure accurate LiDAR point cloud projection under varying image crops. During training, the AdamW optimizer is used with a weight decay of $1\times 10^{-5}$ and an initial learning rate of $1\times 10^{-4}$. The OneCycleLR scheduler is employed for learning rate adjustment. The model is trained for 100 epochs with a batch size of 6, and the number of iterative refinements is set to 4 during training. All the experiments are conducted on a single NVIDIA RTX 4090 GPU.
4.1.3 Evaluation Metrics
@@
Figure 4: Error distributions under four different experimental configurations (Exp1-Exp4).
4.2 Experimental Results
@@
In this section, we analyze the quantitative and qualitative results of our method compared to other methods. Additionally, we examine its runtime performance.
4.2.1 Quantitative Results
@@
As shown in Tab. 2, UniCalib is evaluated against seven state-of-the-art approaches across various initial deviations and datasets, including NetCalib [30], DXQ-Net [10], CALNet [22], PSNet [31], CalibFormer [32], CalibDepth [37] and CalibLG [18]. Fig. 4 illustrates the translation and rotation error distributions from four distinct experiments, confirming that our method consistently achieves high-accuracy calibration results across all tested scenarios.
@@
Exp1: UniCalib achieves a mean translation error of $0.550\mathrm{cm}$ and a mean rotation error of $0.044^{\circ}$. When compared to DXQ-Net, a matching-based method, UniCalib maintains comparable rotation error while greatly reducing the translation error by $28.9\%$. This improvement stems from the PWSF loss, which suppresses unreliable gradients from uncertain regions, leading to more robust depth flow estimation and improved calibration accuracy.
Figure 5: Comparison of LiDAR projection on RGB images under four different experimental configurations (Exp1-Exp4). Each column shows the result with: (a) initial uncalibrated extrinsics, (b) extrinsics estimated by UniCalib, and (c) GT extrinsics.
@@
Exp2&\&3: UniCalib achieves a mean translation error of $0.913\mathrm{cm}$ and a mean rotation error of $0.053^{\circ}$ in Exp2, surpassing CalibFormer, which regresses the extrinsics from a camera image and a sparse LiDAR depth map, by $23.1\%$ and $62.4\%$, respectively. In Exp3, it attains a mean translation error of $2.368\mathrm{cm}$ and a mean rotation error of $0.084^{\circ}$, outperforming CalibLG, another regression-based method, by $37.8\%$ and $65.0\%$. These improvements arise from our intra-modality formulation, which converts inputs to a unified depth representation, thereby ensuring consistent feature extraction for calibration and significantly boosting the accuracy of both translation and rotation estimation.
@@
Exp4: UniCalib achieves a mean translation error of $9.664\mathrm{cm}$ and a mean rotation error of $0.381^{\circ}$ on the KITTI-360 dataset without being trained on it, outperforming CalibDepth, the only open-source method among the baselines, by $48.6\%$ and $84.2\%$, respectively. This strong generalization ability stems from our probabilistic matching within a unified depth representation, ensuring robustness against both visual appearance variations and geometric outliers.
4.2.2 Qualitative Results
@@
As shown in Fig. 5, our estimated extrinsics closely aligns with the ground truth across diverse scenarios and initial deviation conditions, highlighting its robustness to a wide range of initial errors and effectiveness in handling complex environments.
Table 3: Ablation studies on different components in UniCalib
@@
Trans. ($\mathrm{cm}$) $\downarrow$ Rot. ($\mathrm{\SIUnitSymbolDegree}$) $\downarrow$ Trans. ($\mathrm{cm}$) $\downarrow$ Rot. ($\mathrm{\SIUnitSymbolDegree}$) $\downarrow$
@@
    D, U, R, and P denote Completed Dense Depth Map, Unified Feature Encoder, Reliability Map, and PWSF Loss, respectively. “w/o P (L1)” and “w/o P (NLL)” denote replacing PWSF loss with L1 and NLL losses, respectively.
4.2.3 Ablation Study